OpenStudy (karatechopper):

How do you plug in answers to find extraneous solutions from a trig equation?

OpenStudy (xapproachesinfinity):

u just plug the solution like you do with regular equations!? give an example to see it

OpenStudy (karatechopper):

Uhm ok Cosx(tanx)=cosx

OpenStudy (karatechopper):

I got the solutions (degrees) 90, 270, 45, 225

OpenStudy (karatechopper):

Idk how to put in the solutions though to check?

OpenStudy (solomonzelman):

\(\large\color{slate}{ \cos x(\tan x)=\cos x }\) \(\large\color{slate}{ \cos x(\tan x)-\cos x=0 }\) \(\large\color{slate}{ \cos x(\tan x-1)=0 }\) \(\large\color{slate}{ \cos x=0 }\) or \(\large\color{slate}{ \tan x-1=0 }\) \(\large\color{slate}{ \tan x=1 }\)

OpenStudy (solomonzelman):

to check a solution just plug it in for x.

OpenStudy (solomonzelman):

9into the original equation)

OpenStudy (karatechopper):

mhm i know. But then how are 90 and 270 extraneous?

OpenStudy (xapproachesinfinity):

well for instance 90 wouldn't be and acceptable solution since it makes tan undefined

OpenStudy (karatechopper):

I think I may be phrasing myself wrong a bit. Should I put in both cosx=0 and tanx=1 at the same time to check together

OpenStudy (xapproachesinfinity):

any solution that you plug and results in \[\cos x\tan x =\cos x\]

OpenStudy (solomonzelman):

\(\large\color{slate}{ \cos (90)\tan(90)=\cos(90) }\) \(\large\color{slate}{ \cos (270)\tan(270)=\cos(270) }\) tan is sin/cos, and so tan(90) and tan(270) is undefined.

OpenStudy (xapproachesinfinity):

i meant \[\cos x\tan x\ne \cos x\]

OpenStudy (solomonzelman):

saying, that any solution that makes cos(x) be zero, is extraneous, because as you plug it into the tangent, your tangent will be undefined.

OpenStudy (solomonzelman):

xapproachesinfinity, it is not an identity, but there are some solutions.

OpenStudy (xapproachesinfinity):

no i'am saying that any solution that makes the equation not true is extra...

OpenStudy (karatechopper):

Oh I see what you are saying soloman You see, I used to just put in the fractions or whatever cosx equaled.

OpenStudy (karatechopper):

correct xapproach

OpenStudy (xapproachesinfinity):

and 90, 270 are examples

OpenStudy (solomonzelman):

yes, basically, for: \(\large\color{slate}{ \cos (x)\tan(x)=\cos(x) }\) re-writing: \(\large\color{slate}{ \displaystyle \cos (x)\frac{\sin(x)}{\cos(x)}=\cos(x) }\) for \(\large\color{slate}{ \cos (x)\tan(x)=\cos(x) }\) you have 2 sets of solutions: 1) where \(\large\color{slate}{ \cos (x)=0 }\) 2) where \(\large\color{slate}{ \tan (x)=1 }\)

OpenStudy (solomonzelman):

But, considering that, \(\large\color{slate}{ \displaystyle \tan(x)=\frac{\sin(x)}{\cos(x)} }\), any solution that makes cos(x) equal to 0 is extraneous, because tan(x) is undefined.

OpenStudy (xapproachesinfinity):

i think i didn't make my point clear lol thanks solom

OpenStudy (solomonzelman):

oh, i understood:) lol

OpenStudy (solomonzelman):

well, as long the questioner gets it... karatechopper, do you get what I am saying?

OpenStudy (karatechopper):

YEAH I DO! THANKS!